OpenStudy (anonymous):

Find the real-valued solution to the initial value problem 7y''+16y'+16y=0 y(0)=0 y'(0)=5

OpenStudy (anonymous):

First Solve \[ 7 r^2 + 16 r + 16 = 0\\ \begin{array}{c} r=\frac{4}{7} \left(-2-i \sqrt{3}\right) \\ r=\frac{4}{7} \left(-2+i \sqrt{3}\right) \\ \end{array} \] Can you finish it from here?

OpenStudy (anonymous):

First, thank you for replying! and Second, I got up to that and plugged it into the equation and got stuck

OpenStudy (anonymous):

I ended up with y(x)=e^(-8/7x)[C1cos((8/14x(sqrt3))+C2sin((8/14)x(sqrt(3))

OpenStudy (anonymous):

This the general solution \[ y=A e^{-8 x/7} \cos \left(\frac{4 \sqrt{3} x}{7}\right)+B e^{-8 x/7} \sin \left(\frac{4 \sqrt{3} x}{7}\right) \] Find A and B from the initial condition.

OpenStudy (anonymous):

\[ 0 = y(0)= A \] So A =0 Try to get B alone. I will check when I come back form shopping.

OpenStudy (anonymous):

ok thanks so much!

OpenStudy (anonymous):

i solved for y' and ended up getting B=-32.65544457

OpenStudy (anonymous):

and it's sadly incorrect

OpenStudy (anonymous):

\[ y'=-\frac{4}{7} \sqrt{3} A e^{-8 x/7} \sin \left(\frac{4 \sqrt{3} x}{7}\right)-\frac{8}{7} A e^{-8 x/7} \cos \left(\frac{4 \sqrt{3} x}{7}\right)-\frac{8}{7} B e^{-8 x/7} \sin \left(\frac{4 \sqrt{3} x}{7}\right)+\frac{4}{7} \sqrt{3} B e^{-8 x/7} \cos \left(\frac{4 \sqrt{3} x}{7}\right)\\ y'(0)=\frac{4 \sqrt{3} B}{7}-\frac{8 A}{7}=\frac{4 \sqrt{3} B}{7}=5\\ B=\frac{35}{4 \sqrt{3}} \]

OpenStudy (anonymous):

\[ y=\frac{35 e^{-8 x/7} \sin \left(\frac{4 \sqrt{3} x}{7}\right)}{4 \sqrt{3}} \] is your final answer.

OpenStudy (anonymous):

thank you!